1. Sketch the hydrated ions that form when LiBr dissolves in water. Show the inner hydration sphere only.
2. In each of the following pairs of compounds, which compound is likely to be more soluble in CCl4? Explain
a. Br2 or NaBr b. CH3CH2OH or CH3OCH3
c. CS2 or KOH d. I2 or CaF2
3. Polonium has a simple cubic unit cell and an atomic radii of 168 pm. What is the density of polonium in g/cm3?
4. The density of an element is 2.540 g/cm3. The element crystalizes in a face centered cubic lattice and has an atomic radius of 216.37 pm. What element is it?
1
Expert's answer
2020-01-30T04:31:28-0500
  1. CCl4 is a non-polar solvent, so it dissolves non-polar molecules:
  • a) Br2 (non-polar molecule);
  • b) CH3OCH3 (ethers are less polar than alcohols);
  • c) CS2 (KOH is ionic compound, that normally do not dissolve in organic solvents);
  • d) I2 (non-polar molecule).

3.Polonium has a simple cubic unit cell so it has 1 atom per unit cell (1/8 atom at each corner).

Find the mass of one Po atom:

m = 209 g/mol : 6.022 * 1023 atom/mol = 3.470608 * 10-22 g/atom.

For a simple cubic unit the edge length of the unit cell is twice the atomic radius:

X = 168 * 2 = 336 (pm) = 3.36 * 10-8 cm

Find the volume of the cell:

V = (3.36 * 10-8)3 = 3.793 * 10-23 (cm3/atom)

Find the density of Polonium:

D = 3.470608 * 10-22 g/atom : 3.793 * 10-23 cm3/atom = 9.15 g/cm3

Answer: Polonium density is 9.15 g/cm3.

4. For a face centered cubic lattice the edge length of the unit cell is "\sqrt8" times longer than atomic radius:

X = 216.37 * "\sqrt8" = 612 (pm) = 6.12 * 10-8 (cm)

Find the volume of the unit cell:

V = (6.12 * 10-8)3 = 2.292 * 10-22 (cm-3)

Find the mass of atoms in the cell:

m = 2.292 * 10-22 * 2.540 = 5.82168 * 10-22 (g/unit)

As there are 4 atoms in a face centered cubic lattice, find the mass of 1 atom:

m = 5.82168 * 10-22 / 4 = 1.45542 * 10-22 (g/atom)

Find the molar mass of the element:

M = 1.45542 * 10-22 g/atom * 6.022 * 1023 atom/mol = 87.64 g/mol

The molar mass obtained corresponds to Strontium.

Answer: the unknown element is Strontium.